# Training Optimal Large Diffusion Language Models

### v1

@@ 6.1 Scaling Laws
Understanding how scaling affects large language model (LLM) performance has been a central research focus. The seminal work of Kaplan et al. (2020) showed that model performance follows predictable power-law trends with respect to model size, compute, and training data, implying that ever-larger models should yield better results. This paradigm shaped the development of models such as GPT-3 (Brown et al., 2020). However, Hoffmann et al. (2022) challenged this view with Chinchilla, demonstrating that, under a fixed compute budget, optimal performance arises from scaling model size and training data in tandem. This revealed that many prior models, including Gopher (Rae et al., 2021), were undertrained, shifting the field’s understanding toward balanced scaling and more efficient compute utilization.
@@
Subsequent research has refined our understanding of scaling laws, particularly in data-constrained regimes. The Chinchilla laws, though influential, assume effectively unlimited training data. As models scale further, the scarcity of unique, high-quality data has emerged as a critical bottleneck. Muennighoff et al. (2023) introduced a data-constrained scaling law (Equation (7)) to model validation loss under limited data, where repeated exposure reduces the "effective model size" and "effective data size." While this captures diminishing returns, the formulation has a key limitation: it enforces a non-increasing validation loss, whereas in practice repeated epochs inevitably induce overfitting, increasing validation loss due to the bias–variance tradeoff. In this work, we propose a new formulation that addresses this flaw.
@@
Beyond pre-training loss, scaling law research is expanding to downstream task performance (Isik et al., 2024), inference dynamics (Wu et al., 2025b), and theoretical grounding, linking empirical trends to concepts such as data manifold dimensionality (Bahri et al., 2024; Sharma and Kaplan, 2022). This broadening scope underscores the need for more refined laws that integrate model architecture, data quality, and task-specific requirements. Li et al. (2025) further explored hyperparameter scaling, offering practical guidance for pre-training choices.
@@
For DLMs, systematic scaling laws were lacking prior to Quokka. Nie et al. (2024) trained models at low FLOPs budgets to compare scaling trends of AR models and DLMs, marking an important step toward DLM scaling, though their study provided only limited scaling law coefficients & insights.
6.2 Diffusion language models.
@@
Building on the theoretical foundations of DLMs (Lou et al., 2023; Shi et al., 2024; Ou et al., 2024; Sahoo et al., 2024), Nie et al. (2025) trained the first large-scale DLM from scratch, achieving performance competitive with leading open-source AR models (Dubey et al., 2024). In parallel, several commercial DLMs have emerged, demonstrating strong coding and math capabilities while offering significantly lower generation latency (Google DeepMind, 2025; Khanna et al., 2025; Song et al., 2025). Ni and team (2025) further showed that DLMs possess substantially higher data potential than AR models under limited data, enabling so-called "intelligence crossovers" that highlight their advantage in the face of the token crisis (Xue et al., 2023; Muennighoff et al., 2023).
@@
Efforts have also explored hybrid approaches bridging AR and diffusion. Block diffusion (Arriola et al., 2025) performs block-wise diffusion, with block size 1 similar to AR modeling without shift. Dream (Ye et al., 2025) initialized DLMs with AR priors and employed a "shift-by-one" strategy to better retain AR knowledge, offering another effective training paradigm. Recent work has also advanced DLM coders (Gong et al., 2025; Xie et al., 2025), DLM RL scaling (Zhu et al., 2025a), accelerated inference techniques (Wu et al., 2025a), pushing DLMs toward greater practicality and competitiveness.
@@ 7 Discussions
In practice, model training is often constrained by resources beyond compute–leading to deviations from the allocations prescribed by scaling laws. For instance, Llama 3 (Dubey et al., 2024) trained an 8B model with 15T tokens, whereas the Chinchilla law would suggest a 70B model for 2T tokens. Several factors contribute to such deviations: (1) Scaling-optimal allocation is not the only consideration for commercial models; factors such as deployability, customer adoption, and hardware compatibility (e.g., GPU/TPU memory limits) play a decisive role. (2) Compute budgets are not always strict. In many cases, one can effectively "expand" compute by extending training time, making smaller models with more data or epochs more practical than adhering rigidly to scaling predictions. (3) Current compute- and data-constrained scaling laws are limited in scope, and their coefficients can shift across architectures and datasets. Thus, scaling laws should be viewed as high-level guidance on balancing model size, data, and training duration, while precise choices require empirical tuning under specific constraints.
@@ Appendix A Implementation Details
All experiments were conducted with a heavily modified Megatron-LM codebase. Compute-constrained runs and ablations were trained on a subset of the Nemotron-CC corpus (Su et al., 2024), while data-constrained runs used a subset of the c4-en corpus (Raffel et al., 2020). Validation losses were consistently evaluated on the c4-val split, following Muennighoff et al. (2023). Token budgets were randomly sampled from the respective corpora without additional filtering. Model parameters were initialized from a normal distribution with standard deviation 0.02. Architecturally, we adopted a performant configuration combining the GPT-2 tokenizer, RoPE, SwiGLU, pre-layer RMSNorm, bias-free layers, and qk normalization.
@@ B.2 Additive Overfitting Term v2
Table 5: The FLOPs and Tokens allocation predicted by approach 2 and 3. Similar to (Hoffmann et al., 2022), the loss fitting approach under-estimates $N\rightarrow N_{opt}$ for very large models.

### v2

@@ 6.1 Scaling Laws
Understanding how scaling affects large language model (LLM) performance has been a central research focus. The seminal work of kaplan2020scaling showed that model performance follows predictable power-law trends with respect to model size, compute, and training data, implying that ever-larger models should yield better results. This paradigm shaped the development of models such as GPT-3 (brown2020language). However, hoffmann2022training challenged this view with Chinchilla, demonstrating that, under a fixed compute budget, optimal performance arises from scaling model size and training data in tandem. This revealed that many prior models, including Gopher (rae2021scaling), were undertrained, shifting the field’s understanding toward balanced scaling and more efficient compute utilization.
@@
Subsequent research has refined our understanding of scaling laws, particularly in data-constrained regimes. The Chinchilla laws, though influential, assume effectively unlimited training data. As models scale further, the scarcity of unique, high-quality data has emerged as a critical bottleneck. muennighoff2023scaling introduced a data-constrained scaling law (Equation (7)) to model validation loss under limited data, where repeated exposure reduces the "effective model size" and "effective data size." While this captures diminishing returns, the formulation has a key limitation: it enforces a non-increasing validation loss, whereas in practice repeated epochs inevitably induce overfitting, increasing validation loss due to the bias–variance tradeoff. In this work, we propose a new formulation that addresses this flaw.
@@
Beyond pre-training loss, scaling law research is expanding to downstream task performance (isik2024scaling), inference dynamics (wu2025inference), and theoretical grounding, linking empirical trends to concepts such as data manifold dimensionality (bahri2024explaining; sharma2022scaling). This broadening scope underscores the need for more refined laws that integrate model architecture, data quality, and task-specific requirements. li2025predictable further explored hyperparameter scaling, offering practical guidance for pre-training choices.
@@
For DLMs, systematic scaling laws were lacking prior to Quokka. nie2024scaling trained models at low FLOPs budgets to compare scaling trends of AR models and DLMs, marking an important step toward DLM scaling, though their study provided only limited scaling law coefficients & insights.
6.2 Diffusion Language Models
@@
Building on the theoretical foundations of DLMs (lou2023discrete; shi2024simplified; ou2024your; sahoo2024simple), nie2025large trained the first large-scale DLM from scratch, achieving performance competitive with leading open-source AR models (dubey2024llama). In parallel, several commercial DLMs have emerged, demonstrating strong coding and math capabilities while offering significantly lower generation latency (deepmind2025geminiDiffusion; khanna2025mercury; song2025seed). ni2025difflm further showed that DLMs possess substantially higher data potential than AR models under limited data, enabling so-called "intelligence crossovers" that highlight their advantage in the face of the token crisis (xue2023repeat; muennighoff2023scaling).
@@
Efforts have also explored hybrid approaches bridging AR and diffusion. Block diffusion (arriola2025block) performs block-wise diffusion, with block size 1 similar to AR modeling without shift. Dream (ye2025dream) initialized DLMs with AR priors and employed a "shift-by-one" strategy to better retain AR knowledge, offering another effective training paradigm. Recent work has also advanced DLM coders (gong2025diffucoder; xie2025dream), DLM RL scaling (zhu2025llada), accelerated inference techniques (wu2025fast), pushing DLMs toward greater practicality and competitiveness.
@@ 7 Discussions
In practice, model training is often constrained by resources beyond compute–leading to deviations from the allocations prescribed by scaling laws. For instance, Llama 3 (dubey2024llama) trained an 8B model with 15T tokens, whereas the Chinchilla law would suggest a 70B model for 2T tokens. Several factors contribute to such deviations: (1) Scaling-optimal allocation is not the only consideration for commercial models; factors such as deployability, customer adoption, and hardware compatibility (e.g., GPU/TPU memory limits) play a decisive role. (2) Compute budgets are not always strict. In many cases, one can effectively "expand" compute by extending training time, making smaller models with more data or epochs more practical than adhering rigidly to scaling predictions. (3) Current compute- and data-constrained scaling laws are limited in scope, and their coefficients can shift across architectures and datasets. Thus, scaling laws should be viewed as high-level guidance on balancing model size, data, and training duration, while precise choices require empirical tuning under specific constraints.
8 Acknowledgment
We thank Shen Nie, Jiacheng Ye, and Cunxiao Du for their fruitful discussions and pointers.
@@ Appendix A Implementation Details
All experiments were conducted with a heavily modified Megatron-LM codebase. Compute-constrained runs and ablations were trained on a subset of the Nemotron-CC corpus (su2024nemotron), while data-constrained runs used a subset of the c4-en corpus (raffel2020exploring). Validation losses were consistently evaluated on the c4-val split, following muennighoff2023scaling. Token budgets were randomly sampled from the respective corpora without additional filtering. Model parameters were initialized from a normal distribution with standard deviation 0.02. Architecturally, we adopted a performant configuration combining the GPT-2 tokenizer, RoPE, SwiGLU, pre-layer RMSNorm, bias-free layers, and qk normalization.
@@ B.2 Additive Overfitting Term v2
Table 5: The FLOPs and Tokens allocation predicted by approach 2 and 3. Similar to (hoffmann2022training), the loss fitting approach under-estimates $N\rightarrow N_{opt}$ for very large models.